OpenStudy (anonymous):

show that there is an infinite number of solutions to x^2= -1 in the quaternions

ganeshie8 (ganeshie8):

Hey!

OpenStudy (anonymous):

hi

ganeshie8 (ganeshie8):

check the last reply http://math.stackexchange.com/questions/288938/x21-0-uncountable-many-solutions

OpenStudy (anonymous):

Let a,b,c be real numbers satisfying a2+b2+c2=1 and let x=ai+bj+ck. Then x2=(ai+bj+ck)(ai+bj+ck)=a2i2+abij+acik+baji+b2j2+bcjk+caki+cbkj+c2k2=?−(a2+b2+c2)=−1. where are these terms acik+baji+bcjk+caki+cbkj?

OpenStudy (reemii):

\(i^2 = j^2 = k^2 = ijk = -1\). From that (or from the definition..), you see that \(i^2 = ijk = -1\) \(\to i = jk\). Meaning, any product of two letterrs among \(i,j,k\) can be replaced by the third letter, with a + or - sign. https://en.wikipedia.org/wiki/Quaternion#Definition

OpenStudy (kainui):

I've been playing around with this, maybe you guys will be interested. Instead of representing quaternions as i,j,k with a real part, I've found the algebra to be laborious so I have been using this special factorization since ij=k \[q=a+bi+cj+dk =( a+bi) + (c+di)j = u+vj\] so now u and v are complex numbers and they commute with each other, however for any complex number z, we have this interesting property when you commute, it takes the complex conjugate of it. \[zj=jz^*\] So now I'll try to solve \(q^2=-1\) with it. \[q^2 = (u+vj)(u+vj)=u^2+uvj+vju+vjvj \] \[q^2= u^2 + v(u+u^*)j-vv^*\] Now I recognize that this is the real part of u \(\frac{u+u^*}{2} = \Re\{u\}\) and this is the squared length of v(a real number) \(vv^* = |v|^2\). \[q^2=u^2 -|v|^2 +2 v \Re\{u\}j\] So we could sorta cut this off here, and just say \(\Re\{u\}=0\) to get rid of the j part, which would leave just the imaginary part left, make \(u^2 = (i\Im\{u\})^2 = - (\Im\{u\})^2 = - |u|^2\) \[q^2 = -(|u|^2+|v|^2) = -1\] So any 3 numbers we pick that satisfy \(a^2+b^2+c^2=1\) will work (cause the real part of u was thrown out) which happens to be the same result as in @ganeshie8 's reference.

OpenStudy (anonymous):

cool